how can i work this out?
The length of a rectangular playing field is 5 ft less than twice its width.
If the perimeter of the playing field is 230 ft, find the length and width of the field.

Answer by checkley71(8403)

LENGTH=2W-5 & W=WIDTH. THUS
2(2W-5)+2W=230
4W-10+2W=230
6W=230+10
6W=240
W=240/6
W=40 ANSWER FOR THE WIDTH
LENGTH=2*24-5
LENGTH=48-5
LENGTH=43
PROOF
2(2*40-5)+2*40=230
2(80-5)+80=230
2*75+80=230
150+80=230
230=230